OpenStudy (anonymous):

A 945 kg drag race car accelerates from rest to 117 km/h in 0.977 s. What change in momentum does the force produce? Answer in units of kg · m/s

OpenStudy (anonymous):

Well your answer needs to be in KG x m/s Momentum can be written as p=m(mass in kg)*v(m/s) So from this we can say that p=945KG*32.5(m/s)= 3063.125 Kg*m/s Note*: F ▲ t = m ▲ v <-- Memorize this :)